### v1

# Theorem of Cassels and sums of squares of regular functions

Abstract.
Let $C$ be an irreducible nonsingular curve over a formally real field $K$ and let $f$ be a regular function which is a sum of squares in the ring of regular functions of the surface $C\times\mathbb{A}_{K}$. We show that the (sum of squares) length of $f$ is the same in the ring of regular functions of $C\times\mathbb{A}_{K}$ and in the field of rational functions of $C\times\mathbb{A}_{K}$. With this we are able to show that the Pythagoras number of the ring of regular functions is the same as the Pythagoras number of the field of rational functions, for some product algebraic surfaces. This shows that the Pythagoras number of the ring of regular functions of the real plane is 4, which was not known before.
Key words and phrases:
Pythagoras numbers, Waring problem, algebraic surface, quadratic form, sums of squares.
2020 Mathematics Subject Classification:
14P05, 26C99, 11P05
Introduction
Let $f$ be a real valued function defined on a space $X$. Assume that $f$ is a sum of squares of functions on $X$. What is the minimal number of squares needed to represent $f$? In this paper we tackle this problem for regular functions on an algebraic surface, defined over a formally real field $K$. This problem is closely related to that of computing the Pythagoras number of a ring.
Definition 0.1.
For a commutative ring $R$ with identity its Pythagoras number, denoted by $p(R)$, is the smallest positive integer $g$ such that any element which is a sum of squares can be written as a sum of at most $g$ squares. If such number does not exist, we put $p(R)=\infty$.
The problem of computing the Pythagoras number is a generalization of the Waring problem for positive integers. Hence, computing Pythagoras numbers is essentially the Waring problem for squares for more general rings.
For a commutative ring $R$ with identity, determining its Pythagoras number is a rather difficult task. As we will be considering real rings (i.e. rings in which $-1$ is not a sum of squares), let us discuss what is known in this case. It is known that any 1–dimensional affine $\mathbb{R}$–algebra has finite Pythagoras number (see [7]). On the other hand, any affine $K$–algebra over a formally real field $K$ of dimension at least 3 has infinite Pythagoras number. Similarly any regular local ring $R$ of dimension at least 3 with formally real residue field has $p(R)=\infty$ (see [7] for both results).
The case of dimension 2 seems to be the most mysterious one. It is known that $p(\mathbb{R}[[x,y]])=2$ and $p(\mathbb{Z}[[x]])=5$, but $p(\mathbb{Z}[x])=p(\mathbb{R}[x,y])=\infty$ (see [7]). Recently, there has been some development in this area. Pythagoras number of rings of the type $\mathbb{R}[x,y,\sqrt{f(x,y)}]$ where $f(x,y)$ is a polynomial satisfying some mild conditions has been computed in [4]. See also [1] for the $k$-regulous case and [27] for some other problems related to sums of squares. For a 2–dimensional regular local ring $R$ it was shown in [24] that finiteness of the Pythagoras number of the field of fractions of $R$ implies finiteness of the Pythagoras number of $R$ and explicit bound was given.
Computing the Pythagoras number of a number field is rather simple task [20, Chapter 7, Example 1.4]. On the other hand, computing the Pythagoras number of the ring of integers or orders is much more involved, see [9, 11, 16, 17, 18, 26] for some recent developments in this area. It is also worth to mention very recent paper [21], which deals with the Pythagoras number of the field of rational functions of a curve over a number field.
In the literature one can also find treatment of higher Pythagoras numbers (sometimes called Waring numbers) (cf. [2, 10, 14, 15]). Computing Pythagoras number is difficult, and computing higher Pythagoras numbers is much harder, as the theory of higher degree forms is much less developed.
Our main result relates the length of a regular function to the length of this function considered as a rational function. Surprisingly, the Pythagoras number was not known for a ring of regular functions of any surface, not even the real plane. Let us discuss what is known. If $R$ is a ”ring of regular functions” (RFR) in the sense of [19], then any totally positive unit can be written as a sum of squares, and explicit upper bound was given in terms of transcendence degree. For an algebraic surface over a formally real field $K$, the usual ring of regular functions is an RFR, hence the result of Mahe applies. This is however far from computing the Pythagoras number, as there is no control on the lengths of non–units. In a similar spirit one defines the unital Pythagoras number i.e. one considers only units which are sums of squares and not general elements which are sums of squares. It is worth to mention the paper [3, Theorem 5.17] where this notion was studied. The most interesting part is the relation with higher unital Pythagoras numbers. Such relation is known only in very specific cases for usual Pythagoras numbers [15, Conjecture 5.2].
Let us describe the content of the paper. In Section 1 we recall necessary facts from algebraic geometry and theory of quadratic forms needed to prove the main theorem. In Section 2 we provide applications of main result to the problem of Pythagoras numbers. We compute the Pythagoras number of the ring of regular functions on a real plane, which turns out to be 4.
1. Main result
Let us start with preliminary definitions.
Definition 1.1.
We say that a field $K$ is formally real if $-1$ is not a sum of squares of elements of $K$.
In particular, any formally real field $K$ contains $\mathbb{Q}$ and so it is of characteristic 00.
Let $K$ be a formally real field. We will recall all the definitions regarding algebraic geometry and rings of regular functions. In this paper we will avoid the scheme theoretic point of view on algebraic varieties. For the treatment of algebraic sets over a real closed field see [5, Chapter 3]. Let $X\subset K^{n}$ be an algebraic set.
Definition 1.2.
We define the coordinate ring of $X$ as the quotient ring
$K[X]:=K[x_{1},x_{2},\dots,x_{n}]/\mathcal{I}(X)$
where $\mathcal{I}(X)$ is the ideal of all the polynomials vanishing on $X$.
If $X$ is an irreducible algebraic set, then we denote its field of fractions by $K(X)$.
Take any $x_{0}\in X$ and denote by $\mathfrak{m}_{x_{0}}$ the maximal ideal of all polynomial functions vanishing on $x_{0}$.
Definition 1.3.
We say that $x_{0}\in X$ is a nonsingular point if the local ring $K[X]_{\mathfrak{m}_{x_{0}}}$ is a regular local ring.
We say that an algebraic set $X\subset K^{n}$ is nonsingular if every point point is nonsingular.
As we will be interested only in $K$-rational points, we may allow singularities to occur in the algebraic closure of $K$. The crucial object of our study is the following
Definition 1.4.
Let $X\subset K^{n}$ be an algebraic subset. We define the ring of regular functions on $X$ as
$\mathcal{O}(X)=\left\{\,\frac{f}{g}\,|\,f,g,\in K[X],\,g^{-1}(0)=\varnothing\right\}.$
Let us note that similarly as in the scheme theory one can consider sheaves of modules over the sheaf of regular functions. This geometric approach differs substantially from the case of an algebraically closed field (or scheme theoretic point of view). For example, over $\mathbb{R}$, Cartan’s Theorems $A$ and $B$ do not hold, however they do hold after a suitable multiblowup [12, 13].
In algebraic geometry one usually defines the notion of a regular function locally. However, since we deal only with formally real fields and by [5, Proposition 3.2.3] regular functions on $X$ are precisely those that admit a global denominator. In particular, $\mathcal{O}(X)$ is a localization of $K[X]$ at a set consisting of polynomial functions vanishing nowhere on $X$. If $K=\mathbb{R}$ and $X$ is a smooth algebraic set, then we localize at all polynomial functions which does not change sign on $X$. For more general field this is not true, as the following example shows.
Example 1.5.
Let $K=\mathbb{Q}$ and $X=\mathbb{Q}$. Then the function $\frac{1}{x^{2}-2}$ belongs to $\mathcal{O}(X)$ even though it changes sign on $X$.
We may now define the notion of a length, which will be useful in the study of sums of squares.
Definition 1.6.
Let $R$ be a commutative ring with identity. We define the length of an element $a\in R$ to be the smallest positive integer $g$ such that $a$ can be written as a sum of at most $g$ squares.
Example 1.7.
Let $R=\mathbb{Z}$ and $a=7$. In this case it is known that the length of $7$ is $4$.
From the very definition of the length we obtain the following fact
Lemma 1.8.
Let $R$ be an integral domain and $a\in R$. Then the length of $a$ in $R$ is greater than or equal to the length of $a$ in the field of fractions of $R$. ∎
Apart from algebraic geometry, we will use elements of the theory of quadratic forms (see [20]). Let us now recall the Theorem of Cassels [6], where only sums of squares were considered instead of a general quadratic form. The statement we present below can be found in [20, Chapter 1, Theorem 2.2].
Theorem 1.9.
Let $K$ be a field of characteristic different from 2. Let $\varphi$ be a quadratic forms with coefficients in $K$ and $f\in K[x]$ be a polynomial. If $f$ is represented by $\varphi$ over $K(x)$ then it is represented by $\varphi$ already over $K[x]$.
Let $K$ be a formally real field. We will consider product surfaces, hence to avoid awkward formulas such as $K[C\times K]$, we will denote the affine line over $K$ as $\mathbb{A}_{K}$. In this case, the above coordinate ring will be denoted as $K[C\times\mathbb{A}_{K}]$ which turns out to be $K[C][y]$, where $y$ is a single variable.
Lemma 1.10.
Let $C$ be a curve over $K$ and let $x_{0}\in C$ be a nonsingular point. Then the curve $D=\{x_{0}\}\times\mathbb{A}_{K}\subset C\times\mathbb{A}_{K}$ is a zero set of a single polynomial functions. Moreover, this polynomial function can be chosen to be the generator of the prime ideal of all polynomial functions vanishing on $D$.
Proof.
The coordinate ring of the surface $C\times\mathbb{A}_{K}$ is equal to $K[C][y]$ for a single variable $y$. One can find a function $g\in K[C]$ such that $g$ generates the maximal ideal $\mathfrak{m}_{x_{0}}$ of the point $x_{0}$. Then the function $g$ vanishes precisely on $\{x_{0}\}\times\mathbb{A}_{K}$. Additionally, $g$ is the generator of the prime ideal of polynomial functions vanishing on $D$.
We are now ready to prove the main result.
Theorem 1.11.
Let $K$ be a formally real field. Let $C$ be an irreducible nonsingular curve over $K$, and $X=C\times\mathbb{A}_{K}$ be a product surface. Let $f\in\mathcal{O}(X)$ be a regular function which is a sum of squares. Then the lengths of $f$ in $K(X)$ and $\mathcal{O}(X)$ are equal.
Proof.
Let $f\in\mathcal{O}(X)$ be a sum of squares. By Lemma 1.8 the length of $f$ in $\mathcal{O}(X)$ is greater than or equal to the length of $f$ in $K(X)$.
By the previous lemma, the curve of the form $\{x_{0}\}\times\mathbb{A}_{K}$ is cut out globally by a single equation for any $x_{0}\in C$. We may write $f=f_{1}h^{2}$, where the zero set of $h$ consists precisely of finite number of curves of the aforementioned type. Replacing $f$ by $f_{1}$ we may assume that $f$ does not not vanish on any curve of the form $\{x_{0}\}\times\mathbb{A}_{K}$ for any $x_{0}\in C$.
Of course, $K[C][y]\subset K(C)[y]$. Denote by $N$ the length of $f$ in $K(X)$ and let $\varphi=\sum_{i=1}^{N}x_{i}^{2}$ be a quadratic form. By the very definition of $N$, $\varphi$ represents $f$ over $K(X)$, hence, by the Theorem of Cassels also over $K(C)[y]$. Let $f=\sum_{i=1}^{N}\frac{f_{i}^{2}}{g^{2}}$ be a presentation such that $f_{i}\in K[X]$ and $g\in K[C]$. Any zero $x_{0}\in C$ of $g$ corresponds to a curve of the form $\{x_{0}\}\times\mathbb{A}_{K}$ in $C\times\mathbb{A}_{K}$ and those are given by a single equation by Lemma 1.10. Hence, we can factor them from $g$ and all $f_{i}$ for $i=1,2,\dots,N$. Consequently, $g$ does not vanish and so we obtain representation of $f$ as a sum of squares of regular functions of length $N$. ∎
2. Applications and further remarks.
In this section we will apply the above result to the theory of Pythagoras numbers. We will also discuss possible generalization of the main result to the case of more general rings as well as to higher even powers instead of squares.
2.1. Pythagoras numbers
Corollary 2.1.
With the assumptions of Theorem 1.11 we obtain $p(\mathcal{O}(X))=p(K(X))$.
Proof.
It is enough to note that if $f\in K(X)$ is a rational function of length $N$, then one has a presentation $f=\frac{\sum_{i=1}^{N}f_{i}^{2}}{g^{2}}$ for $f_{i},g\in K[X]$. Necessarily, the lengths of $f$ and $fg^{2}$ are equal in $K(X)$ and additionally $fg^{2}\in K[X]$. By Theorem 1.11, the lengths of $fg^{2}$ in $\mathcal{O}(X)$ and $K(X)$ are equal, hence the thesis.
The above corollary reduces the problem of finding the Pythagoras number of the ring of regular functions to that of the field of rational functions. In some very specific cases this is known. We are able to deduce the following.
Corollary 2.2.
For $X=\mathbb{R}^{2}$ we have $p(\mathcal{O}(\mathbb{R}^{2}))=4$.
Results of Scheiderer [25] imply the following.
Theorem 2.3.
Let $p\in\mathbb{R}[x,y]$ be a nonnegative polynomial. Then, for sufficiently large $N$ the polynomial $(x^{2}+y^{2}+1)^{N}p$ is a sum of squares of polynomials.
However, no bound on the number of squares is known in this case. By the proof of Theorem 1.11 and Corollary 2.2 we able to refine this result.
Corollary 2.4.
Let $p\in\mathbb{R}[x,y]$ be a nonnegative polynomial. Then, there exists a polynomial $h\in\mathbb{R}[x]$ such that $h^{2}p$ is a sum of 4 squares of polynomials.
We get the precise number of summands at the cost of not knowing the precise denominator, which depends only on 1 variable.
Let $R$ be a regular local ring of dimension 2. Denote by $F$ the field of fractions of $R$. Problem 5 in [7] asked about a relation between $p(R)$ and $p(F)$. Scheiderer [24] proved that $p(R)\leq 4p(F)-4$. In particular, $p(R)$ is finite if and only if $p(F)$ is finite. We are able to strengthen this result, at least for some regular local rings of geometric origin.
Corollary 2.5.
With the assumptions of Theorem 1.11, let $x_{0}\in X$ be any nonsingular point. Let $\mathcal{O}(X)_{\mathfrak{m}_{x_{0}}}$ be the regular ring of the point $x_{0}$. Then $p(\mathcal{O}(X)_{\mathfrak{m}_{x_{0}}})=p(K(X))$.
Proof.
Since $\mathcal{O}(X)_{\mathfrak{m}_{x_{0}}}$ is a localization of $\mathcal{O}(X)$ we have a following chain of inequalities
$p(\mathcal{O}(X))\geq p(\mathcal{O}(X)_{\mathfrak{m}_{x_{0}}})\geq p(K(X)).$
By the previous corollary, the first and the last quantities are equal, hence the result. ∎
As a consequence, we formulate the following conjecture.
Conjecture 2.6.
Let $R$ be a regular local ring of dimension 2 and $K$ be its field of fractions. Then $p(R)=p(K)$.
Of course, $p(R)\geq p(K)$, hence this conjecture is only interesting if $p(K)$ is finite.
2.2. Lack of Theorem of Cassels for coordinate rings of curves.
It would be valuable to generalize Theorem 1.11 to the rings of regular functions of products of two irreducible nonsingular curves. This would easily follow from Theorem of Cassels for coordinate rings of curves. However, such a generalization is false as the following example shows. Let $K=\mathbb{R}$ and $k\in\mathbb{R}[x]$ be a nonlinear polynomial of odd degree in a single variable $x$. Denote by $X=\{y^{2}-k(x)=0\}\subset\mathbb{R}^{2}$ a planar curve.
Proposition 2.7.
With the above notation:
  • a)
    The Pythagoras number of $\mathbb{R}[X]$ is 3 or 4.
  • b)
    The Pythagoras number of the field $\mathbb{R}(X)$ is 2.
Proof.
Part $a)$ is precisely [7, Proposition 3.8]. For the part $b)$, observe that $\mathbb{R}(X)$ has transcendence degree 1 over $\mathbb{R}$, hence by the result of Pfister [20, Corollary 3.4] we have $p(\mathbb{R}(X))\leq 2$. Straightforward calculations shows that $1+\bar{x}^{2}$ is not a square in $\mathbb{R}(X)$, where $\bar{x}$ is the image of $x$ in $\mathbb{R}[X]$.
With the above at hand, let $\varphi=x_{1}^{2}+x_{2}^{2}$ be a quadratic form. There exists an element $f\in\mathbb{R}[X]$ which is a sum of at least three squares but not of two squares. On the other hand $\varphi$ represents this element in the fraction field. Hence, Cassels Theorem cannot be generalized to coordinate rings of curves, not even nonsingular ones.
2.3. Lack of Theorem of Cassels for higher powers.
Let $n>1$ be a positive integer. Theorem of Cassels also does not hold for sums of higher powers instead of squares. This follows from the fact that there are polynomials in a single variable $x$ which are sums of $2n$-th powers of rational functions, but no sums of $2n$-th powers of polynomials [22]. Let us discuss the case of fourth powers, as examples for higher powers can be produced similarly.
Consider the family of polynomials $f_{m}=x^{4}+mx^{2}+1$ [23, Exercise 7.4.4]. An easy application of Cauchy-Schwartz-Bunyakovsky inequality shows that $f_{m}$ is a sum of fourth powers of polynomials only for $m\in[0,6]$. On the other hand, by the Theorem of Becker (see [2, 8]) $f_{m}$ is a sum of fourth powers of rational functions for $m\in[0,+\infty)$.
Acknowledgment
The author is very grateful to Bartłomiej Bychawski and Piotr Miska for their comments, remarks and many fruitful discussions.